A girl whose mass is 40kg walks up a flight of 20steps each 150mm height in 10 seconds. Find the average power developed. (g = 10 metre per second square)




1
Expert's answer
2021-04-06T13:53:07-0400

"P=mgh/t=40\cdot 10\cdot 20\cdot 0.15/10=120\ (J)"